OpenStudy (anonymous):

Three roses will be selected for a flower vase. The florist has 1 red, 1 white, 1 yellow, 1 orange, 1 pink from which to choose. How many different three rose selections can be formed from the 5 roses? Can someone explain it well cause I'm confused.

OpenStudy (anonymous):

5 C 3 , i guess !

OpenStudy (anonymous):

what is the probability that 3 roses selected at random will contain 1 red, 1 white, and 1 pink?

OpenStudy (anonymous):

We want three, and there are 5. Let's label them R=red, W=white, Y=yellow, O=orange, P=pink. The different combinations we can have are: RWY RWO RWP RYO RYP RPO WYO WYP WOP YOP Since the order doesn't matter, that is, RWY = RYW = YWR, that's 10 different combinations, which, as mm said, is 5C3. What that function means is actually 5!/3!*2! = 5*4*3*2*1/3*2*2*1 = 120/12

OpenStudy (anonymous):

There is only one combination that produces RWP, therefore, the probability of this being the random selection is the number of ways this can happen (ie, 1) divided by the total possible outcomes (ie, 10). Therefore, the answer is 1/10 = .1

OpenStudy (anonymous):

how would you do that with a npr and ncr though?

OpenStudy (anonymous):

5=n 3=r, hence 5C3. The permutation (nPr) would indicate that RWY<>RYW<>YRW, which I don't believe is true for this problem (unless the instructions indicate otherwise).

OpenStudy (anonymous):

Does that make sense?